1. A specific product has demand during lead time of 350 units, with a standard deviation during lead time of 80 units. Management wants a 99% service level. a) What value of Z should be applied? b) How many units should be carried as safety stock? c) What is the appropriate reorder point?
Answer #1

1.a)

z value should be normsinv(0.99) = 2.326347874 for 99% service level

b)

Safety stock = z*standard deviation during lead time = 2.326347874*80 = 186.1078299 = 186 units

c)

Reorder point = Demand during lead time + safety stock = 350+186 = 536 units